OpenStudy (anonymous):

If the Original height, h, of a triangle is doubled, and the original of base, b, is tripled, which expression would represent the area of the resulting triangle? a, ½‎bh b, 3bh c, 6bh d, 12bh

OpenStudy (anonymous):

3bh it'll be b

OpenStudy (anonymous):

Thank you

OpenStudy (primeralph):

you're not allowed to say answers

OpenStudy (anonymous):

And on top of that, his answer is wrong. @Martha.Ln do you know the formula to find the area of a triangle?

OpenStudy (anonymous):

Not exactly

OpenStudy (anonymous):

Okay...this gets sticky...but I'm gonna go there anyway. The formula is 1/2*B*H That formula holds true regardless of how you twist or rearrange a triangle's measurements. In this case, your question asked "which expression would represent the area" of a triangle. So which of your answers "represents the area" of a triangle?

OpenStudy (anonymous):

So, it's ½‎bh, a ?

OpenStudy (anonymous):

Correct!!

OpenStudy (anonymous):

^^ Thanks alot!

OpenStudy (anonymous):

No problem! Glad I could help! :-) BTW, just to illustrate your problem further, Suppose you had a triangle with a height of 2 and a base of 2 Its area would be 1/2*B*H = 1/2*2*2 = 2 Now double its height and triple its base, and you have a triangle with a height of 4 and a base of 6 The area of this one is STILL 1/2*B*H = 1/2*6*4 = 12 Taadaaa!! :-) I'll go away now... :-)